# Theorem Prover as a Judge for Synthetic Data Generation

### v1

@@ 6 Conclusion
While TP-as-a-Judge demonstrates strong capabilities in verifying intermediate reasoning for synthetic data generation, several limitations persist. Firstly, theorem prover verification is highly effective in mathematical reasoning but remains challenging in other domains, making its extension beyond mathematics an open research question. While formalisation enhances verification compared to standard methods, it is not flawless, requiring further refinement for broader applicability. Secondly, our current implementation primarily focuses on algebra, counting, probability, and problem-solving datasets. Expanding to additional mathematical areas could improve generalisation and model performance. Thirdly, dataset complexity is limited by the LLM’s ability to reliably solve problems. When generating more challenging problems, the model often fails, restricting the complexity of synthetic data and preventing the creation of more advanced mathematical questions. Next, computational constraints limit scalability, particularly for models exceeding 8 billion parameters, requiring substantial resources for large-scale verification and training. Finally, the iterative nature of formalisation incurs a high computational cost, as multiple refinement cycles are needed to ensure correct execution. Addressing these challenges would enable more efficient and scalable synthetic data generation.
@@ Appendix A Detailed Breakdown of Dataset
The dataset breakdown is presented in Table 7. It comprises 2,562 samples for SFT and 473 pairs (a total of 946 samples) for DPO training, resulting in an overall count of 3,508 samples.
@@ A.1 SIMQA dataset Creation Details
In creating our datasets, we use OpenAI GPT-4o models, which includes gpt-4o-2024-08-06 and gpt-4o-2024-11-20 and ChatGPT-4o-latest alternatively for the SIMQA synthetic dataset creation process. For API details regarding GPT-4o, please refer to https://platform.openai.com/docs/models/gpt-4o
@@ A.3 Experimental Settings
LLM-Generated Questions: All questions are generated using a few-shot setting. The specific few-shot examples are provided in Section G.
@@ Appendix B Details on hyperparameters Tuning
For RLTPF, we use a two-stage training strategy consisting of one epoch of Supervised Fine-Tuning (SFT) followed by three epochs of Direct Preference Optimisation (DPO) for reinforcement learning. Our hyperparameter selection was configured in a similar way with DPO Rafailov et al. (2024), with using the default seed (seed=42), and a slight configuration in the learning rate, as shown below:
@@
The initial fine-tuning phase utilises a batch size of 4 per device, with gradient accumulation over 16 steps, effectively increasing the batch size to 64. We adopt a linear learning rate scheduler with an initial learning rate of and a warmup phase of 150 steps. To improve training stability, we apply gradient clipping with a maximum norm of 0.3 and enable gradient checkpointing to reduce memory overhead. The optimisation strategy employs RMSprop, and we save model checkpoints at the end of each epoch.
@@ Appendix C Consistency Check
We conducted a consistency check by evaluating three runs of llama-3.2-3b-instruct and llama-3.1-8b-instruct using different random seeds (seed 42, seed 2103 and seed 1234). The results, summarised in Tables 8 and 9, indicate that the average accuracy remains stable across runs, with relatively low standard deviation. However, GSM8K-Symbolic and AIME exhibit higher standard deviation, suggesting that the true accuracy may be higher than the reported values.
@@ Appendix D Evaluation Metrics
Similar to Leang et al. (2024), majority of our datasets use exact match as our evaluation metrics. For the OlympiadBench datasets, where exact match is not a suitable evaluation metric, we used the GPT-4o-mini model as a benchmark to assess how closely the answers of the model aligned with the ground truth. To prevent the model from generating reasoning that could influence its decision, we input only the final three sentences, which typically contain both the answer and the correct solution. The task was to determine whether these two elements matched, without providing any intermediate reasoning. The prompt is shown below:
@@ Appendix F Dataset Details
Appendix G Few-shot examples for Synthetic Data Creation
In our experiments, we employed the Reverse Question-Answering method, which utilises only the solution as the few-shot example during synthetic data generation. This approach ensures that the model learns to reconstruct problem statements solely from the provided solutions.
@@ Appendix G Few-shot examples for Synthetic Data Creation
Appendix H Manual Annotation Details
We conduct a manual annotation process to verify the correctness of the theorem prover’s validations. Manual annotation is performed by a mathematics and statistics bachelor’s graduate, who is specifically instructed to solve the mathematical problems. To ensure consistency in manual verification, we randomly sampled 54 instances from the GSM8K-synthetic dataset (see examples in Appendix I).

### v2

@@ 6 Conclusion
While TP-as-a-Judge demonstrates strong capabilities in verifying intermediate reasoning for synthetic data generation, it has several limitations. Firstly, theorem prover verification is highly effective in mathematical reasoning but remains challenging in other domains, making its extension beyond mathematics an open research question. While formalisation enhances verification compared to standard methods, it is not flawless, requiring further refinement for broader applicability. Secondly, our current implementation primarily focuses on algebra, counting, probability, and problem-solving datasets. Expanding to additional mathematical areas could improve generalisation and model performance. Thirdly, dataset complexity is limited by the LLM’s ability to reliably solve problems. When generating more challenging problems, the model often fails, restricting the complexity of synthetic data and preventing the creation of more advanced mathematical questions. Next, computational constraints limit scalability, particularly for models exceeding 8 billion parameters, requiring substantial resources for large-scale verification and training. Finally, the iterative nature of formalisation incurs a high computational cost, as multiple refinement cycles are needed to ensure correct execution. Addressing these challenges would enable more efficient and scalable synthetic data generation.
Acknowledgements
We thank the reviewers for their useful feedback and comments. We also thank Pasquale Minervini, Yifu Qiu, Aryo Pradipta Gema, Xuanli He, Christina Du, Rohit Saxena, Zhaowei Wang, Isaac Ong and Dongwei Jiang for their constructive feedback and insights. We appreciate the computational resources provided by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. We thank netmind.AI for an additional contribution of computational resources.
@@ Appendix A Detailed Breakdown of Dataset
The dataset breakdown is presented in Table 8. It comprises 2,562 samples for SFT and 473 pairs (a total of 946 samples) for DPO training, resulting in an overall count of 3,508 samples. Based on the datasets generated, Theorem Prover as a Judge is generalisable to various domains within mathematics, and iterative autoformalisation has increased its generalszability across domains, covering areas such as arithmetic, number theory, algebra, and others. However, in domains beyond mathematics, iterative autoformalisation may be ineffective. Additional analysis of each component’s of Theorem Prover as a Judge contribution to dataset creation can be referred to Appendix 11.
@@ A.1 SIMQA dataset Creation Details
In creating our datasets, we use OpenAI gpt-4o models, which include gpt-4o-2024-08-06 and gpt-4o-2024-11-20 and ChatGPT-4o-latest alternatively for the SIMQA synthetic dataset creation process. For API details regarding gpt-4o, please refer to https://platform.openai.com/docs/models/gpt-4o.
@@ A.3 Experimental Settings
LLM-Generated Questions: All questions are generated using a few-shot setting. The specific few-shot examples are provided in Section H.
@@ Appendix B Details on hyperparameters Tuning
For RLTPF, we use a two-stage training strategy consisting of one epoch of Supervised Fine-Tuning (SFT) followed by three epochs of Direct Preference Optimisation (DPO) for reinforcement learning. Our hyperparameter selection was configured in a similar way with DPO Rafailov et al. (2023), using the default seed (seed=42), and a slight configuration in the learning rate, as shown below:
@@
The initial fine-tuning phase uses a batch size of 4 per device, with gradient accumulation over 16 steps, effectively increasing the batch size to 64. We adopt a linear learning rate scheduler with an initial learning rate of and a warmup phase of 150 steps. To improve training stability, we apply gradient clipping with a maximum norm of 0.3 and enable gradient checkpointing to reduce memory overhead. The optimisation strategy uses RMSprop, and we save model checkpoints at the end of each epoch.
@@ Appendix C Consistency Check
We conducted a consistency check by evaluating three runs of llama-3.2-3b-instruct and llama-3.1-8b-instruct using different random seeds (seed 42, seed 2103 and seed 1234). The results, summarised in Tables 9 and 10, indicate that the average accuracy remains stable across runs, with relatively low standard deviation. However, GSM8K-Symbolic and AIME exhibit higher standard deviation, suggesting that the true accuracy may be higher than the reported values.
@@ Appendix D Evaluation Metrics
Similar to Leang et al. (2024), the majority of our datasets use exact match as our evaluation metrics. For the OlympiadBench datasets, where exact match is not a suitable evaluation metric, we used the gpt-4o-mini model as a benchmark to assess how closely the answers of the model aligned with the ground truth. To prevent the model from generating reasoning that could influence its decision, we input only the final three sentences, which typically contain both the answer and the correct solution. The task was to determine whether these two elements matched, without providing any intermediate reasoning. The prompt is shown below:
@@ Appendix F Dataset Details
Appendix G Breakdown of Theorem Prover as a Judge
| Category | Subcategory | Error (%) |
| Question Formalisation | Autoformalisation Error Alignment Check Error | 13.10 10.00 |
| Answer Formalisation | — | 5.40 |
Autoformalisation presents inherent challenges due to the difficulty of ensuring that autoformalised questions align with their original counterparts. Most recent works involve multiple calls to LLMs and expensive automated theorem proving techniques Li et al. (2024b); Poiroux et al. (2024). Similarly, our generated pipeline is crucial at each step, as components either facilitate correct formalisation (e.g., CoMAT during question formalisation) or ensure verification accuracy (e.g., alignment checks).
The decomposition into question and answer formalisation is based on existing theorem provers, as research demonstrates that question formalisation improves answer verification. Each aspect of our theorem prover-as-judge approach has been carefully curated to enhance autoformalisation and data quality.
Autoformalisation is an inherently challenging problems due to the difficulty of ensuring that autoformalised questions align with their original counterparts. Our generated pipeline is crucial at each step because components either facilitate correct formalisation (e.g., CoMAT Leang et al. (2024) during question formalisation) or ensure verification accuracy (e.g., alignment check). The decomposition into question and answer formalisation is based on existing theorem provers Jiang et al. (2023b), as research shows question formalisation improves answer verification. Each aspect of theorem prover as judge has been carefully curated to enhance autoformalisation and data quality.
Within the question-formalisation phase , the CoMAT step, improves performance by approximately 20%, while our autoformalisation component, enhances answer formalisation. Auto-Informalisation, and alignment checks serve as metrics to verify that CoMAT and autoformalisation are performed correctly. This additional layer is essential as there are no definitive metrics to validate question formalisation correctness. Misalignments potentially caused by errors may result in disposal before proceeding to answer formalisation. Removing any component could compromise data quality (eg. omitting CoMAT may reduce autoformalisation by 20%, while eliminating alignment checks might allow incorrect but seemingly verified autoformalisation to progress to answer formalisation, affecting the verification of model reasoning.
As shown in Figure 4, 23.10% of total samples resulted in question formalisation error, 5.4% in answer formalisation error, while the remaining 71.50% were categorised as verified or false. The detailed breakdown of filtering components is shown in Table 11. While we have not analysed CoMAT, it could potentially be evaluated using alignment check as well. These statistics provide clear analysis of each component.
Appendix H Few-shot Examples for Synthetic Data Creation
In our experiments, we used the Reverse Question-Answering method, which uses only the solution as the few-shot example during synthetic data generation. This approach ensures that the model learns to reconstruct problem statements solely from the provided solutions.
@@ Appendix H Few-shot Examples for Synthetic Data Creation
Appendix I Manual Annotation Details
We conduct a manual annotation process to verify the correctness of the theorem prover’s validations. Manual annotation is performed by a mathematics and statistics bachelor’s graduate, who is specifically instructed to solve the mathematical problems. To ensure consistency in manual verification, we randomly sampled 54 instances from the GSM8K-synthetic dataset (see examples in Appendix J).
Few Shot Examples for GSM8K
Few Shot Examples for MATH
Algebra
Counting and Probability
Appendix J Examples for Manual Annotation
Appendix K Example of Iterative Autoformalisation
In the process of Theorem Prover as a Judge, there are two main components: question formalisation and answer formalisation. During question formalisation, we follow four steps (denoted ): applying the CoMAT method, autoformalisation, Auto-Informalisation, and Alignment Check. We used the example from the previous appendix (Manual Example 6.2) as an illustration. In this example, we present how autoformalisation works during question formalisation and then proceed to answer formalisation, providing a detailed analysis of how iterative autoformalisation refines the Lean code into a correct version over three iterations. When one is formalised as False (6.1) and the other as Verified (6.2), the model will be trained using Direct Preference Optimisation. The examples and prompts have been modified to remove unicode,